COMPUTER and express in scientific notation
1. 540.89 +46.435+33.8-13.89
2. 1.372x0.026782x0.004843/0.79003
3.50.5+825.819- (4.167 +21.56)
4.8.189/(0.06022 x 0.04483x0.8865)
1
Expert's answer
2020-11-22T11:15:28-0500

1. 540.89 +46.435+33.8-13.89 = 607.235 = 6.07235 x 102


2. 1.372x0.026782x0.004843/0.79003

=2.1867 x 10-4


3.50.5+825.819- (4.167 +21.56)= 850.592 =8.50592 x 102


4.8.189/(0.06022 x 0.04483x0.8865)= 3421.7056 = 3.4217056 x 103